# Gravitational-wave memory effects in the Damour-Esposito-Farèse extension of Brans-Dicke theory

### v1

@@ Abstract
Gravitational-wave memory effects are lasting changes in the strain and its time integrals. They can be computed in asymptotically flat spacetimes using the conservation and evolution equations in the Bondi-Sachs framework. Modified theories of gravity have additional degrees of freedom with their own asymptotic evolution equations; these additional fields can produce differences in the memory effects in these theories from those in general relativity. In this work, we study a scalar-tensor theory of gravity known as the Damour-Esposito-Farèse extension of Brans-Dicke theory. We use the Bondi-Sachs framework to compute the field equations in Bondi-Sachs form, the asymptotically flat solutions, and the leading gravitational-wave memory effects. Although Damour-Esposito-Farèse theory has additional nonlinearities not present in Brans-Dicke theory, these nonlinearities are subleading effects; thus, the two theories share many similarities in the leading (and some subleading) solutions to hypersurface equations, asymptotic symmetries, and types of memory effects. The conservation equations for the mass and angular momentum aspects differ between the two theories, primarily because of the differences in the evolution equation for the scalar field. This leads to differences in the time dependence of the gravitational-wave memory signals that are produced during the quasicircular inspiral of compact binaries. These differences, however, are of second-order in a small coupling parameter of these theories, which suggests that it would be challenging to use memory effects to distinguish between these two theories.
@@ I Introduction
What is now referred to as the displacement GW memory effect was first studied in the 1970s Zel’dovich and Polnarev (1974) (see also Smarr (1977); Turner (1978); Epstein (1978); Bontz and Price (1979) and earlier related discussion in Newman and Penrose (1966)). The study of GW memory effects has expanded since their first computation several decades ago largely because of developments in their theoretical study and observational search prospects.
@@
The observational prospects for detecting the displacement memory had been studied prior to the observation of GWs Braginsky and Grishchuk (1985); Braginsky and Thorne (1987); Kennefick (1994); Favata (2009a); Pollney and Reisswig (2011), but these studies focused on individual sources of GWs. After the detection of binary-black-hole (BBH) mergers by LIGO, it was realized that detecting evidence for the memory effect in the entire population of BBH mergers was possible Lasky et al. (2016). Subsequent forecasts based on the BBH population inferred from the LVK detections have indicated that the memory effect in the BBH population is likely to be observed during the fifth LVK observing run Hübner et al. (2020); Boersma et al. (2020); Grant and Nichols (2023). There are active searches for the memory effect in BBH populations using LVK data Hübner et al. (2020, 2021); Cheung et al. (2024) and from individual events from subsolar mass BBHs Ebersold and Tiwari (2020) or with pulsar timing arrays Wang et al. (2015); Agazie et al. (2024). It will also be possible to detect the memory with future GW detectors on the ground Johnson et al. (2019); Grant and Nichols (2023); Goncharov et al. (2024) and in space Favata (2009a); Islo et al. (2019); Goncharov et al. (2024); Gasparotto et al. (2023); Inchauspé et al. (2024).
@@
There have been developments in the study of memory effects from a theoretical perspective, too. First, GW memory effects were found to be part of an infrared triangle (see, e.g., Strominger and Zhiboedov (2016); Strominger (2017)) which relates them to the asymptotic symmetries of the spacetime and to soft theorems of gravitational scattering Weinberg (1965). Second, several additional GW memory observables were predicted that generalize the displacement memory. The displacement memory is named as such because it can be measured by the change in separation of two comoving freely falling observers after the passage of a burst of GWs (i.e., their geodesic deviation). If the observers have a relative velocity, then there will be a residual change in their displacement after the burst of waves that depends on their initial velocity and the time integral of the shear. This subleading GW memory Flanagan et al. (2019); Grant and Nichols (2022) (also called drift memory Grant (2024)) is closely connected with the spin Pasterski et al. (2016); Flanagan and Nichols (2017); Nichols (2017) and center-of-mass (CM) Nichols (2018) GW memory effects. Observers with relative acceleration can measure a hierarchy of higher GW memory effects, which depend on the initial time derivatives of the relative acceleration and two or more integrals of the GW strain Flanagan et al. (2019); Grant and Nichols (2022); Grant (2024); Grant and Mitman (2024); Siddhant et al. (2024).
@@
All of these GW memory effects can be computed from (i) “fluxes” (or “pseudofluxes”), which are nonlinear in the Bondi metric functions and vanish in the absence of radiation, and (ii) “charges,” which can be constructed to be constant in the absence of radiation. In some cases, the charges and fluxes are associated with underlying symmetries of asymptotically flat spacetimes. This is shown most definitively for the displacement memory effect, which is closely connected to the supertranslation symmetries of asymptotically flat spacetime in the Bondi-Metzner-Sachs (BMS) group Bondi et al. (1962); Sachs (1962) via the infrared triangle discussed above. The spin and CM memory effects are related to the extended Barnich and Troessaert (2010a, b, 2011) or generalized Campiglia and Laddha (2014, 2015) BMS algebras, though not quite as straightforwardly as with the standard BMS group Compère et al. (2018). The charges from which the higher memory effects also may be related to a $w_{1+\infty}$ algebraic structure Strominger (2021); Freidel et al. (2022, 2024); Compère et al. (2022); Geiller (2024). However, in this paper we will restrict to computations of the displacement and spin memory effects sourced by the nonlinear Christodoulou (1991); Blanchet and Damour (1992); Flanagan and Nichols (2017); Nichols (2017) (or flux) terms.
We focus on the leading displacement and spin memory effects in this paper, because our focus will be on memory effects in gravitational theories beyond GR. GWs from compact-binary sources have been studied extensively in these theories (see, e.g., Yunes and Siemens (2013); Yunes et al. (2024)), and they have been constrained by the LVK observations Abbott et al. (2021a, b). Memory effects in beyond-GR theories, however, have been studied in less detail, but there have been a handful of calculations performed in the past few years. Brans-Dicke (BD) theory Tahura et al. (2020, 2021); Hou and Zhu (2020a, b) and dynamical Chern-Simons gravity Hou et al. (2022a, b) were studied using the Bondi-Sachs framework. The displacement GW memory effect also was computed in most general scalar-vector-tensor theory Heisenberg et al. (2023, 2024), which itself encompasses a wide class of specific beyond-GR theories, using the Isaacson effective stress-energy tensor in beyond-GR theories, instead. There have also been linearized-theory calculations of asymptotic symmetries and their connections to GW memory effects in the Lorentz-violating Einstein-Æther theory Hou et al. (2024) and in diffeomorphism-invariant theories containing tensor degrees of freedom that closely resemble those in GR Hou (2024).
@@
In both GR and beyond GR theories, computing memory effects can be performed perturbatively, approximately using balance laws, or numerically in full generality. In GR, memory effects have been computed during the inspiral phase of a binary system using the flux terms and post-Newtonian (PN) Favata (2009b, 2011); Ebersold et al. (2019) or self-force Elhashash and Nichols (2024); Cunningham et al. (2024) waveforms. These calculations are consistent with full PN calculation at 4PN order that do not make use of the balance laws Trestini and Blanchet (2023); Blanchet et al. (2023a, b) when compared at the overlapping PN orders for each calculation. Post-Newtonian calculations using the balance laws have also been performed in BD theory Tahura et al. (2021). The fluxes in the Bondi-Sachs framework also can be used to compute memory effects approximately using full, numerical-relativity (NR) waveforms Nichols (2017); Mitman et al. (2021, 2020); Khera et al. (2021); Grant and Mitman (2024), because NR simulations that extract waveforms using extrapolation often do not capture GW memory effects (see, e.g., Boyle et al. (2019)).
@@
Although NR simulations using extraction do not capture the displacement memory effect, there are two closely related numerical frameworks that do. The first is Cauchy-characteristic evolution (CCE), which evolves NR data with a finite outer boundary to future null infinity after the finite-domain NR simulation has been run, and Cauchy-characteristic matching (CCM), which performs both evolutions coupled and simultaneously (see, e.g., the review Winicour (2009)). CCE and CCM compute gravitational waveforms at future null infinity that include memory effects in a particular BMS frame (see the review Mitman et al. (2024) for more details about the choice of the frame). The NR code SpECTRE Moxon et al. (2020, 2023); Ma et al. (2024a); Deppe et al. (2024), for example, is under development and is implementing both the CCE and CCM algorithms.
@@
While most of the development of CCE and CCM has been performed in the context of GR, there has been recent work in Ma et al. (2024b), which solved the Einstein-Klein-Gordon system (by incorporating a massless scalar field into the SpECTRE CCE). This will be an important step towards computing GW memory effects from compact binary coalescences in non-GR theories, which often have additional scalar degrees of freedom. More specifically, the work of Ma et al. (2024b) could be applied to Bergmann–Wagoner scalar-tensor (ST) theories Wagoner (1970); Bergmann (1968) or Damour-Esposito-Farèse (DEF) theory with a single scalar field Damour and Esposito-Farese (1992), in the Einstein frame. NR simulations of black-hole–neutron-star systems in DEF theory were performed, which could be used as input to CCE (see, e.g., Ma et al. (2023)).
Useful cross-checks of the GW memory effects computed from these anticipated CCE and CCM simulations in ST theories will be having expressions for the conservation equations that the memory effects must satisfy and having analytical PN results of GW memory effects. Our focus in this paper will be to derive such balance laws and compute the PN memory waveforms in DEF theory with a single scalar field in Bondi-Sachs coordinates, which are used by the CCE and CCM frameworks. We will focus on the displacement and spin memory effects, as they are the leading electric and magnetic-parity memory effects that have been computed in GR (see, e.g., Nichols (2017); Mitman et al. (2021); Grant and Mitman (2024); Siddhant et al. (2024)) as well as in BD theory Tahura et al. (2020, 2021).
@@
DEF theory is a useful example of a ST theory to consider because it contains additional nonlinearities that produce nontrivial strong-field dynamics (such as spontaneous scalarization for neutron stars Damour and Esposito-Farese (1993), and it also reduces to BD theory when the coupling function in the action between gravity and the scalar field becomes a constant. The Bondi-Sachs framework is well adapted for studying nonlinear dynamics of massless fields in the wave zone, and it will allow us to understand what strong-field features of DEF theory also have an impact on the propagation of tensor and scalar GWs far from the source. Because it reduces to BD theory in a particular limit, we can also more easily determine which features of the memory waveforms are associated with having an additional, noninteracting scalar degree of freedom, and which are associated with the additional nonlinearities in the dynamics of the scalar field.
@@
In this paper, we will focus on the similarities and differences of aspects of asymptotically flat spacetimes and memory effects in DEF and BD theories. We will write the results in a way that highlights the similarities with the results of Refs. Tahura et al. (2020, 2021) when they occur and makes manifest the differences between the two theories. Most of these differences arise from the evolution of the scalar field and, in particular, the subleading dynamical part of the field. We will also discuss how these differences ultimately enter into the PN inspiral memory waveforms that are computed from compact-binary sources on quasicircular orbits.

### v2

@@ Abstract
Gravitational-wave memory effects are lasting changes in the strain and its time integrals. They can be computed in asymptotically flat spacetimes using the conservation and evolution equations in the Bondi-Sachs framework. Modified theories of gravity have additional degrees of freedom with their own asymptotic evolution equations; these additional fields can produce differences in the memory effects in these theories from those in general relativity. In this work, we study a scalar-tensor theory of gravity known as the Damour-Esposito-Farèse extension of Brans-Dicke theory. We use the Bondi-Sachs framework to compute the field equations in Bondi-Sachs form, the asymptotically flat solutions, and the leading gravitational-wave memory effects. Although Damour-Esposito-Farèse theory has additional nonlinearities not present in Brans-Dicke theory, these nonlinearities are subleading effects in an expansion in inverse luminosity distance; thus, the two theories share many similarities in the leading (and some subleading) solutions to hypersurface equations, asymptotic symmetries, and types of memory effects. The conservation equations for the mass and angular momentum aspects differ between the two theories, primarily because of the differences in the evolution equation for the scalar field. When used to compute the time dependence of the gravitational-wave memory signals from quasicircular inspirals of binary neutron stars and black-hole–neutron-star binaries, they give rise to new features that are of second-order in a small coupling parameter in this theory. The observational bound that the coupling is small suggests that it would be challenging to use memory effects during the inspiral to distinguish between Brans-Dicke and Damour-Esposito-Farèse theories. Nevertheless, these results can be used to analyze and interpret memory effects from numerical-relativity simulations of binaries in this theory.
@@ I.1 Background on memory effects in GR
The GW memory effect was first predicted in the 1970s Zel’dovich and Polnarev (1974) (see also Smarr (1977); Turner (1978); Epstein (1978); Bontz and Price (1979) and earlier related discussion in Newman and Penrose (1966)) and was computed in the context of linearized gravity. It has subsequently been called “linear GW memory,” because Christodoulou Christodoulou (1991) predicted a nonlinear memory effect, and Blanchet and Damour Blanchet and Damour (1992) independently computed the effect in the context of multipolar post-Minkowski expansion (see, e.g., Blanchet (2024) for a review of the post-Minkowski formalism). This latter effect has been called “nonlinear GW memory.” The study of GW memory effects has expanded since these first computations several decades ago largely because of developments in their theoretical study and observational prospects, which we review next.
@@
The measurement prospects for detecting the displacement memory had been studied prior to the observation of GWs Braginsky and Grishchuk (1985); Braginsky and Thorne (1987); Kennefick (1994); Favata (2009a); Pollney and Reisswig (2011), but these studies focused on individual sources of GWs, including binary black hole (BBH), black-hole neutron-star (BH-NS) and binary neutron star (BNS) mergers. After the LVK’s detection of many BBH mergers, it was realized that detecting evidence for the memory effect in the entire population of BBH mergers would be possible Lasky et al. (2016). Subsequent forecasts based on the BBH population inferred from the LVK detections have indicated that evidence for the memory effect in the BBH population is likely to be measured during the fifth LVK observing run Hübner et al. (2020); Boersma et al. (2020); Grant and Nichols (2023). There are active searches for the memory effect in BBH populations using LVK data Hübner et al. (2020, 2021); Cheung et al. (2024) and from individual events from subsolar mass BBHs Ebersold and Tiwari (2020) or with pulsar timing arrays Wang et al. (2015); Agazie et al. (2024). It will also be possible to detect the memory with future GW detectors on the ground Johnson et al. (2019); Grant and Nichols (2023); Goncharov et al. (2024) and in space Favata (2009a); Islo et al. (2019); Goncharov et al. (2024); Gasparotto et al. (2023); Inchauspé et al. (2025) from individual BBH systems. As more BH-NS mergers are detected by the LVK in upcoming observing runs and by next-generation detectors, the memory effect can be detected from these systems and used to distinguish BBH from BH-NS mergers Tiwari et al. (2021).
There have been developments in the study of GW memory effects from a theoretical perspective, too. First, GW memory effects were found to be part of an infrared triangle (see, e.g., Strominger and Zhiboedov (2016); Strominger (2017)) which relates them to the asymptotic symmetries of the spacetime and to soft theorems of gravitational scattering Weinberg (1965). Second, several additional GW memory observables were predicted that generalize the linear and nonlinear effects of Zel’dovich and Polnarev (1974); Christodoulou (1991), respectively. To describe these many new effects and distinguish them from the memory effects of Zel’dovich and Polnarev (1974); Christodoulou (1991), this led to new nomenclature for the effects, which we now review.
@@
First, the linear and nonlinear effects of Zel’dovich and Polnarev (1974); Christodoulou (1991) are collectively referred to as “displacement memory” Nichols (2017). It was given this name because it can be measured by the change in separation of two comoving freely falling observers after the passage of a burst of GWs (i.e., the solution of the geodesic deviation that depends on initial relative separation) Flanagan et al. (2019). However, if the observers have a relative velocity, then there will be a residual change in their displacement after the burst of waves that depends on their initial velocity and the time integral of the shear Flanagan et al. (2019); this effect is independent of the displacement memory Grant (2024). This subleading GW memory Flanagan et al. (2019); Grant and Nichols (2022) (also called “drift memory” Grant (2024)) is closely connected with the spin Pasterski et al. (2016); Flanagan and Nichols (2017); Nichols (2017) and center-of-mass (CM) Nichols (2018) GW memory effects, which have also been related to infrared triangles Strominger (2017) (and quadrilaterals Compère et al. (2018)). Observers with a relative acceleration can measure a hierarchy of “higher GW memory effects,” which depend on the initial time derivatives of the relative acceleration and two or more integrals of the GW strain Flanagan et al. (2019); Grant and Nichols (2022); Grant (2024); Grant and Mitman (2024); Siddhant et al. (2024).
@@
Analogously to the linear and nonlinear contributions to the displacement memory, all of these higher GW memory effects can be expressed in terms of (i) integrals of “fluxes” (or “pseudofluxes”), which are nonlinear in the metric variables and vanish in the absence of radiation, and (ii) changes in “charges,” which can be constructed to be constant in the absence of radiation (and which include some terms that are linear in the metric variables) Grant and Nichols (2022). In some cases, the charges and fluxes are associated with underlying symmetries of asymptotically flat spacetimes. This has been shown most comprehensively for the displacement memory effect, which is closely connected to the supertranslation symmetries of asymptotically flat spacetime in the Bondi-Metzner-Sachs (BMS) group Bondi et al. (1962); Sachs (1962) (and the conjugate supermomentum charges) via the infrared triangle discussed above. The spin and CM memory effects are related to the extended Barnich and Troessaert (2010a, b, 2011) or generalized Campiglia and Laddha (2014, 2015) BMS algebras through the charges and fluxes, though not quite as straightforwardly as with the standard BMS group Compère et al. (2018). The charges from which the higher GW memory effects arise also may be related to a (loop) $w_{1+\infty}$ algebraic structure, though this is still being investigated (see, e.g., Strominger (2021); Freidel et al. (2022, 2024); Compère et al. (2022); Geiller (2025)).
@@
There are several methods to compute GW memory effects, including using different perturbative techniques, the charge and flux relationships described above, and certain numerical-relativity (NR) simulations. Memory effects have been computed during the inspiral phase of a binary system using the (nonlinear) flux terms and post-Newtonian (PN) Favata (2009b, 2011); Ebersold et al. (2019) or self-force Elhashash and Nichols (2025); Cunningham et al. (2024) waveforms. These calculations are consistent with full PN calculations at 4PN order which do not make use of the balance laws Trestini and Blanchet (2023); Blanchet et al. (2023a, b) for the common PN orders in each calculation method. The (nonlinear) fluxes in the Bondi-Sachs framework also can be used to compute memory effects approximately using NR waveforms Nichols (2017); Mitman et al. (2021, 2020); Khera et al. (2021); Grant and Mitman (2024). This is necessary, because NR simulations that extract waveforms from the computational domain using extrapolation do not capture all the GW memory effects (see, e.g., Boyle et al. (2019)).
@@
Although NR simulations using extraction do not resolve the displacement memory effect, there are two closely related numerical frameworks that do. The first is Cauchy-characteristic evolution (CCE), which evolves NR data at a finite outer boundary to future null infinity (after the finite-domain NR simulation has been run). The second is Cauchy-characteristic matching (CCM), which evolves on a domain of finite extent and a compactified infinite domain simultaneously (see, e.g., the review Winicour (2009)). CCE and CCM compute gravitational waveforms at future null infinity that include memory effects in a particular BMS frame (see the review Mitman et al. (2024) for more details about the choice of the frame). The NR code SpECTRE Moxon et al. (2020, 2023); Ma et al. (2024); Deppe et al. (2024), for example, is under development and is implementing both the CCE and CCM algorithms.
I.2 Memory effects in gravitational theories beyond GR
Computing memory effects in beyond-GR theories is a newer enterprise. The oscillatory GWs from compact-binary sources have been studied extensively in a number of beyond-GR theories (see, e.g., Yunes and Siemens (2013); Yunes et al. (2024)), and they have been constrained by the LVK observations Abbott et al. (2021a, b). Memory effects in beyond-GR theories, however, have been studied in less detail. There have been a handful of calculations performed in the past few years, including in Brans-Dicke (BD) theory Tahura et al. (2021a, b); Hou and Zhu (2021a, b) and dynamical Chern-Simons gravity Hou et al. (2022a, b), which were studied using the Bondi-Sachs framework. Post-Newtonian calculations of the displacement and spin memory have been performed in BD theory Tahura et al. (2021b) for compact binary inspirals. [^1] Note that black holes in many scalar-tensor theories satisfy no-hair theorems (see, e.g., Capuano et al. (2023)), so BBH systems are not expected to be sources of scalar waves or scalar memory effects. Thus, the calculations in Tahura et al. (2021b) were most relevant for BH-NS and BNS systems. The displacement GW memory effect also was computed in a general scalar-vector-tensor theory Heisenberg et al. (2023, 2024) (which encompasses a wide class of specific beyond-GR theories) using the Isaacson effective stress-energy tensor Isaacson (1968a, b) in beyond-GR theories (see also, e.g., Stein and Yunes (2011)). There have been linearized-theory calculations of asymptotic symmetries and their connections to GW memory effects in the Lorentz-violating Einstein-Æther theory Hou et al. (2024) and in diffeomorphism-invariant theories containing tensor degrees of freedom that closely resemble those in GR Hou (2024).
One obstacle to computing memory effects in beyond GR theories has been the limited number of systematic analyses of the asymptotic field equations, which identify an appropriate space of asymptotically flat solutions and conservation or evolution equations that can be used to compute memory effects (as in GR). These analytical analyses are also helpful for identifying the asymptotic symmetries, conserved quantities, and other elements of the infrared triangles (e.g., Seraj (2021)). A second limitation is the restricted number of NR simulations of compact-binary mergers in beyond GR theories (a few examples of such simulations in scalar-tensor theories include Barausse et al. (2013); Shibata et al. (2014); Palenzuela et al. (2014)). To use the asymptotic field equations to compute memory effects, NR waveforms that lack the memory can be input into these equations to obtain the unresolved memory signal that is missing from the NR simulations. Finally, it would be beneficial to have beyond-GR NR simulations that use CCE or CCM to analyze and compare with the conservation-law calculations.
@@ I.2 Memory effects in gravitational theories beyond GR
Progress is being made on the NR simulations with and without CCE or CCM. Specifically, NR simulations of BH-NS binaries in Damour-Esposito-Farèse (DEF) theory with a single scalar field Damour and Esposito-Farese (1992) were performed recently. These simulations started early enough in the inspiral that they could be compared with PN calculations (see Ma et al. (2023)). In addition, while most of the development of CCE and CCM has been performed in the context of GR, there has been recent work in Ma et al. (2025), which solved the Einstein-Klein-Gordon system by incorporating a massless scalar field into the framework. This will help with computations of GW memory effects from compact binary coalescences in scalar-tensor (ST) theories of the Bergmann–Wagoner types Wagoner (1970); Bergmann (1968) or in DEF theory, in which the field equations take the form of the Einstein-Klein-Gordon system in a particular conformal frame.
I.3 Objectives and results of this paper
To further enable the calculation and interpretation of GW memory effects in beyond-GR theories, it will be important to continue developing the systematic analyses of the asymptotic field equations and to perform PN calculations of the GW memory effects to match with the inspiral portion of the NR simulations. Our focus in this paper will be to derive the conservation laws and compute the PN memory waveforms in DEF theory with a single scalar field in Bondi-Sachs coordinates. We focus on DEF theory specifically because of the recent high-accuracy simulations in DEF theory Ma et al. (2023) and because it is a representative of general ST theories with one scalar field. We use Bondi-Sachs coordinates, because they are well adapted to the structure of asymptotically flat spacetimes and to the coordinates used in the CCE and CCM frameworks.
@@ I.3 Objectives and results of this paper
In addition, DEF theory contains a rich phenomenology, because of nonlinear interactions of the scalar field with itself, which produce nontrivial strong-field dynamics (such as spontaneous scalarization for neutron stars Damour and Esposito-Farese (1993)). These nonlinearities, in principle, could allow for stronger GW memory effects in all polarizations of the waves. The Bondi-Sachs framework is well adapted for studying nonlinear dynamics of massless fields in the wave zone, and it will allow us to understand what strong-field features of DEF theory also have an imprint on the generation and propagation of GWs far from the source. Another feature of DEF theory that will prove useful is that it reduces to BD theory when the coupling function in the action between gravity and the scalar field becomes a constant. This will allow us to more easily compare with our prior work Tahura et al. (2021a, b), which we henceforth refer to as Papers I and II, respectively. In particular, we can more easily determine which features of the memory waveforms are associated with having an additional, noninteracting scalar degree of freedom as in BD theory, and which are associated with the additional nonlinearities in the dynamics of the scalar field as in DEF theory.
@@
Thus, in this paper, we will focus on the similarities and differences of aspects of asymptotically flat spacetimes and memory effects in DEF and BD theories, and we will write the results in a way that highlights the similarities and differences with the results of Papers I and II. For the memory effects, we will focus on the displacement and spin memory effects, because they are, respectively, the largest amplitude electric and magnetic-parity memory effects from compact binary sources that have been computed in GR (see, e.g., Nichols (2017); Mitman et al. (2021); Grant and Mitman (2024); Siddhant et al. (2024)) as well as in BD theory (Papers I and II). We find that most differences in the field equations and memory effects arise from the evolution of the scalar field, especially the subleading (in an expansion in the inverse Bondi radial coordinate) dynamical part of the scalar field. We also compute how these differences affect the PN inspiral memory waveforms that are computed from compact-binary sources (with at least one NS) on quasicircular orbits.
We now summarize our main findings. In our calculations of the DEF field equations in Bondi-Sachs coordinates, we found that the evolution and conservation equations, when expanded in inverse Bondi-coordinate radius, can be written in a form identical to those in BD theory; however, to write them as such, the metric functions in the Bondi expansion and the subleading part of the scalar field must be modified by nonlinear terms in the leading dynamical part of the scalar field. This result has several implications. For computing the flux (nonlinear) contributions to the displacement and spin memory effects, the formal expressions used will be the same as those in BD theory (see Papers I and II). However, because the radiative data that enters into these calculations for a given source in the two theories will not be the same, the time dependence of the memory signals will generically differ.
As an example of this different time dependence, we discuss the calculation of the PN memory signal during the quasicircular inspiral of a compact binary, which is most relevant for BH-NS or BNS systems. There we find that there would be difference in the Newtonian-order memory signals in BD and DEF theories. However, the differences arise at quadratic order in a coupling parameter that is of order $10^{-5}$. Thus, we expect that differences in the inspiral memory signals between BD and DEF theory (at least during the inspiral) would be too small for current or future GW detectors to measure. We do not perform a detection study of the memory in DEF theory using the PN waveforms because, as in GR, the amount of signal-to-noise ratio (SNR) accumulated during the inspiral is a small fraction of the total SNR, which comes about primarily during the merger and ringdown or disruption (see, e.g., Lasky et al. (2016); Tiwari et al. (2021)). Given that there are not enough full NR waveforms in DEF theory available, we do not perform such a detection study either. However, our results in this paper will allow us to perform such studies once the NR waveforms are available.
I.4 Organization and conventions of this paper